HUBBLE SPACE TELESCOPE DAILY REPORT #1783 PERIOD COVERED: 0000Z (UTC) 12/19/96 - 0000Z (UTC) 12/20/96 Daily Status Report as of 355/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The observations were completed as planned, and no problems were = reported. 1.2 Completed WF/PC-2 6194 (WF/PC-2 Cycle 5: Polarization Proposal) The WF/PC-2 was used to make polarization observations of star = BD+64D106 at four different spacecraft roll angles. This should permit a = detailed and independent calibration of each polarizer quad. The = observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Two Sets of WF/PC-2 6363 (The Environments and Host = Galaxies of BL-Lac Objects) The WF/PC-2 was used to make snapshot survey observations of = BL-Lac objects QSO-0754+1 in order to study their environment and their = host galaxies. The goal is to investigate systematic trends in host = galaxy type, unusual morphological features, nearby companion galaxies, = etc. The observations were executed as scheduled, and no anomalies were = noted. 1.4 Completed GHRS/2 6722 (Probing the Chromospheric Structure of = Alpha Tau) The High Resolution Spectrograph (Side Two) was used to make = spectrographic observations of the K5 III class star Alpha Tau (aka, = HD29139 or Aldeberan). The presence of fluorescence H II and Ca II lines = provides an opportunity to probe the structure of the chromosphere. A = wavelength calibration image was also taken. The observations were = completed as planned, and no problems were reported. 1.5 Completed WF/PC-2 6909 (WF/PC-2 Cycle 6 Calibration: Earth Flats) The WF/PC-2 was used to make observations of the bright Earth to = obtain streak flats which are used to construct high quality flat fields = for the WF/PC-2 filter set. The observations were executed as scheduled, = and no anomalies were noted. 1.6 Completed FOS/RD 6912 (FOS Cycle 6: Dark Monitoring) The Faint Object Spectrograph (Red Side) was used to make = measurements of the internal background. All exposures were internals, = and the spectrographic observations were made in the image mode. The = observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Two Sets of FOS/RD 6914 (FOS Cycle 6: Location of = Spectra and Wavelength Calibration - Red Side) The Faint Object Spectrograph (Red Side) was used to make both a = TALED and a wavelength calibration image. All observations were made = using internal light sources. This proposal will be executed a total of = 12 time during cycle 6 (i.e., once per month). The observations were = executed nominally, with no anomalies. 1.8 Completed FOS/BL 6682 (Geometry and Generalizability of the = Reflected Light Model for Seyfert Galaxies) The Faint Object Spectrograph (Blue Side) was used to make = spectropolarimetry observations of the Seyfert (class 2) galaxy = ESO381-G08 in order to study its geometry and the generalizablity of its = reflected light model. The goal is to determine if the polarization = components of all Seyfert class two galaxies have Seyfert class one = characteristics. Approximately 40 minutes into this two-and-a-half hour = proposal, a loss of lock occcurred. The observations were completed as = scheduled, with no reported effect on science. 1.9 Completed FOS/BL 6539 (Ultraviolet Spectroscope of Seyfert 2 = Nuclei) The Faint Object Spectrograph (Blue Side) was used to make = ultraviolet spectrographic observations of the nucleus of the Seyfert = class two galaxy NGC-5135. The objective is to determine whether or not = the featureless continuum is produced by hot massive stars in a = circumnuclear starburst region. A wavelength calibration image was also = taken. The observations were completed as planned, and no problems were = reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 8 Successful : 8 A loss of lock occurred at354/1939Z. HSTAR 5852 was written. 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Operations Notes: HST operations personnel transferred to the SAMS backup facility = at 355/0730Z in preparation for the SM2 JIS. 3.0 SIGNIFICANT FORTHCOMING EVENTS: JIS #7 (Rendezvous) will be conducted today from approximately = 9:00 a.m. until approximately 9:00 p.m. /DMH